OpenStudy (anonymous):

State whether each of the following is an odd function, an even function or neither an even nor an off function a) G(x)=x/x^4 - 4 b) k(x)=3x -|x|

hartnn (hartnn):

Find G(-x) by replacing x in G(x) by -x. If G(-x) = G(x), then the function is even If G(-x) = -G(x), then the function is odd else neither. same for K(x)

hartnn (hartnn):

could u find G(-x) ?

OpenStudy (anonymous):

No, I need help finding G(-x)

hartnn (hartnn):

wherever u se 'x' in G(x), replace it by -x.

OpenStudy (anonymous):

I'm even more confused. Please simplify it for me

OpenStudy (anonymous):

for odd function f(x)=-f(-x) for even function f(x)=f(-x)

OpenStudy (calculusfunctions):

That is exactly the same thing @hartnn said.

OpenStudy (anonymous):

Could please help me solve at least one so i may use it as a guide?

OpenStudy (anonymous):

@calculusfunctions @shunyao I guess my question i dont know how to solve it, pls

OpenStudy (calculusfunctions):

First replace every x in the equation with -x. What do you get?

OpenStudy (anonymous):

G(-X)=-x/x^4-4 and K(-x)= -3x+|x|

OpenStudy (calculusfunctions):

@odehye that's not clear! Answer my question is the equation I wrote correct or not?

OpenStudy (calculusfunctions):

@shunyao giving out answers is not helpful. Teach, explain, but don't just write the solutions. Please delete that!

OpenStudy (anonymous):

@calculusfunctions not helpful ?

OpenStudy (calculusfunctions):

@shunyao yes!

OpenStudy (anonymous):

@calculusfunctions i deleted it, you may show us your correct steps

OpenStudy (calculusfunctions):

First @odehye I need you to answer me. Is the first equation\[G(x)=\frac{ x }{ x ^{4}-4 }\]the correct equation?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

OK so then Step 1: replace x with negative x. In other words, G(-x) = ?

OpenStudy (calculusfunctions):

If you don't know then don't be ashamed to just tell me.

OpenStudy (anonymous):

I dont

OpenStudy (anonymous):

@calculusfunctions replace x with negative x that is just what i did

OpenStudy (calculusfunctions):

OK then I appreciate your honesty. So suppose we have\[F(x)=\frac{ x -4 }{ x ^{2}-2x+16}\]}\]then \[F(-x)=\frac{ -x -4 }{ (-x)^{2}-2(-x)+16 }\]Do you see what I did? Can you follow this example to now tell me what G(-x) would be?

OpenStudy (calculusfunctions):

@shunyao I know. I didn't say your answer was incorrect. I just said you should be explaining and thus guiding the person asking the question to the correct solution.

OpenStudy (calculusfunctions):

@odehye you still don't understand?

OpenStudy (anonymous):

G(-x)=-x/-x^4-4

OpenStudy (calculusfunctions):

@odehye do you not see my example and how the brackets are used?

OpenStudy (calculusfunctions):

Please try again!

OpenStudy (calculusfunctions):

For example, in my example the next step would be\[F(-x)=\frac{ -x -4 }{ x ^{2}+2x +16 }\]Do you see?

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

OK then once again G(-x) = ?

OpenStudy (anonymous):

@odehye